OpenStudy (anonymous):

Divide 35b5 + 20ab3 + 20a2b2 by 5b2. What is the quotient?

OpenStudy (callisto):

\[\frac{35b^5 + 20ab^3 + 20a^2b^2}{5b^2}\]\[=\frac{35b^5}{5b^2} + \frac{ 20ab^3}{5b^2} + \frac{ 20a^2b^2}{5b^2}\]=cancel the common factor of each numerator and denominator.